OpenStudy (anonymous):

a body is thrown vertically up with a speed of 10 m/s from the top of a wall of height 3.45m the height from which another body should be dropped simultaneously such that two bodies reach the ground in same time (g=10m/s2)

OpenStudy (dumbcow):

using g = 10 solve the quadratic \[-5t^{2}+10t+3.45 = 0\] t = 2.3 Then use \[-5t^{2}+h = 0\] plug in value for t and solve for h h = 26.45

OpenStudy (anonymous):

ha(t) = (g*t^2)/2 + va*t + ha g = -10 va = 10 ha = 3.45 t = ? 0 = -5t^2 + 10*t + 3.45 t = 2.3 hb(t) = (gt^2)/2 + vb*t + hb g = -10 vb = 0 hb = ? t = 2.3 0 = -5(2.3^2) + hb hb = (aprox)26.45m

OpenStudy (vincent-lyon.fr):

@shameer : why do you post this question THREE TIMES? It was already widely discussed and solved in the link below. You are wasting people's time, working on and answering an already solved problem. http://openstudy.com/users/shameer1#/updates/4fa62981e4b029e9dc35e6e7

OpenStudy (anonymous):

the answer is 24.5m can anybody solve for it

OpenStudy (vincent-lyon.fr):

We proved earlier that this answer is NOT correct! Are you not convinced?

OpenStudy (anonymous):

nobody got this aanswer correct

OpenStudy (vincent-lyon.fr):

People who write books and problems sometimes do mistakes too, you know! Stop worrying with this problem and go on! Either something is wrong in the way it is written, or there was a mistake in the answer provided.

OpenStudy (anonymous):

i think u r right thanxxxx

OpenStudy (vincent-lyon.fr):

Even the grammar in this problem sucks.

OpenStudy (anonymous):

and sorry toooo!!

OpenStudy (vincent-lyon.fr):

ok! :) Trust OS for wise answers!